Suppose Micron technology sold today an issue of bonds with a 15-year maturity, a $1000 par value, a 10 percent annual coupon, and semi-annual interest payments. The bonds are callable six years after they are issued. If the bonds were called, Micron Technology would pay a call premium of 10 percent and six months extra interest.

a) Two years after the bonds were issued, the going rate of interest on bonds such as these fell to 8 percent. At what price would these bonds sell at the end of year 2?

b) What is the yield to first call 2 years after the initial offering if the conditions of part(a) hold?

(Show all work please)
Answer #1

a. Bond Price after 2 years = "=PV(rate,nper,pmt,fv)" "=PV(0.04,26,50,1000)" = $1159.83

Rate = Annual Yield / 2 = 8% / 2 = 4%

nper = Number of periods = 13 years * 2 periods = 26 periods

pmt = Interest = $1000 * 5% = $50

FV = $1000

b. yield to first call =RATE(nper,pmt,pv,fv)" = RATE(8,50,-1159.83,1150) = 4.22%

Annual YTC = Semi Annual YTC * 2 = 4.22% *2 = 8.44%

nper = Number of periods = 4 years * 2 periods = 8 periods

pmt = Interest = $1000 * 5% = $50

FV = $1000 + 100 + 50 = $1150